Question  2

Evaluate the definite integral         


"\intop"10 18e3x+1dx (round to an integer).


 

1.      322

2.      311

3.      932

4.      965


1
Expert's answer
2021-07-14T18:13:48-0400

To solve this integral, we have to consider that


"e^{3x+1}=e^u
\\ u=3x+1, u(1)=4, u(0)=1
\\ \implies du=3dx"


With that information we proceed and find:


"I=∫^1_0 18e^{3x+1}dx=6(∫^1_0 3e^{3x+1}dx)
\\ \implies I=6(∫^4_1 e^{u}du)=6 \left( \Big \lbrack e^{u}\Big \rbrack_1^4 \right)=6(e^4-e^1)"


"\implies I\approxeq 6(54.5981-2.7183)\approxeq311.2788\approxeq311"


In conclusion, "\intop"10 18e3x+1dx (round to an integer) is equal to option 2. 311.


Reference:

  • Varberg, D. E., Purcell, E. J., & Rigdon, S. E. (2007). Calculus with differential equations. Pearson/Prentice Hall.